### v1

@@ 2.3.1 Identification of noise in biased models
In general, defining the bias term as a GP process requires the selection of suitable kernels. Typical choices are Radial Basis Functions (RBF) or Matérn kernels, aimed at regression problems. In the implementation of the GP, a prescribed white noise of variance $\sigma_{n}^{2}$ is usually added as in Equation 7. It provides numerical stability and, in some cases, represents the measurement error $\epsilon_{\text{meas}}$. Due to its flexible structure as a GP, the bias term will absorb $\epsilon_{\text{meas}}$, thus including it explicitly is not required. It is generally not possible to distinguish between $\epsilon_{\text{model}}$ and the non-prescribed $\epsilon_{\text{meas}}$. We will then consider any variation in the model discrepancy as part of $\epsilon_{\text{model}}$ and susceptible to be explained by the bias term. If more than one distinct output observation is available per input, the fitted GP tends to solve the regression problem and will not capture that variability. We consider these variations in the observations as noise to be identified, hence a noise-aware kernel with a non-vanishing variance at the observation points is required in those cases.
Homoscedastic and heteroscedastic noise kernels are available to be added to the GP. Homoscedastic ones such as white noise kernels allow us to identify constant noise added to the observations. As the bias term aims to compensate for the discrepancy in the output of the model and $\epsilon_{\text{meas}}$ is assumed to be prescribed, homoscedastic kernels in the bias GP would capture additional observation noise of non-physical nature, such as deficient data collection. A typical homoscedastic kernel is a white noise kernel with noise variance $\sigma_{n}^{2}$ such that
@@
where $g(x_{i})$ represents the evaluation of the regressor built on the pairs $(\bm{\xi}_{i},\sigma_{i}^{2})$. In the implementation used in this work45, $g(x_{i})$ is obtained from applying kernel regression with the training points as anchor points $\hat{\bm{\xi}}$. This kernel can be added analogously to the homoscedastic one to the regressor kernel of the bias GP, where its contribution to the covariance matrix $C_{HN}$ is presented as a diagonal matrix with different entries obtained from the kernel evaluation:
@@
The introduction of noise-aware kernels in the bias allows quantifying the variability in the observations without modifying the deterministic computational model. These kernels can be implemented in both KOH and OGP approaches without needing to further modify the methodology presented in the previous sections. In the case of OGPs, the orthogonality condition is not affected by the noise introduction, as it is part of the kernel evaluation.
@@ 2.3.2 Model enhancement through bias’ input domain extension
One of the most prominent advantages of implementing the bias term independently from the model is the possibility of including information in the bias that enriches and guides the model inference without needing to modify the computational model itself. This is an aspect that, to the knowledge of the authors, has not been explored in the context of structural models, and it has the potential to enhance the predictions generated by simulation-based digital twins. Being $f(x,\theta)$ the computational model, we can extend the model with a bias $b(x,\eta)$ such that $\eta$ are parameter coordinates not considered in $f$ but for which observations are available. Therefore, obviating the measurement noise $\epsilon_{\text{meas}}$, the modified Equation 1 for the final model would be
@@
which would be defined over the extended domain $(x,\eta)\in X\times H$. For example, a temperature-agnostic model defined over the spatial domain $X$ could be extended to consider the temperature by including it in the bias term. This approach requires the datasets used for calibration to be labelled for the extension variable as well as for the original ones. Such is the case for many digital twins, especially for large structures such as bridges, where abundant information from varied sources is gathered independently for monitoring purposes but not necessarily included in the model definition.
The introduction of an extension of the domain requires minor modifications to the inference workflow. In particular, the input training data is composed of pairs of $(x,\eta)$, while the computational model is only defined for $x$. Nevertheless, the GP for the bias takes as output training data the residuals, which are defined over $(x,\eta)$, which allows for keeping the associated pairs of $x$ and $\eta$. Therefore, for the training data of the GP, we define as input dataset the pairs $(\hat{\mathbf{x}},\hat{\bm{\eta}})$ and as output dataset the residuals for a given sampled $\theta$
$\hat{\mathbf{r}}(\hat{\mathbf{x}},\hat{\bm{\eta}},\theta)=y(\hat{\mathbf{x}},\hat{\bm{\eta}})-f(\hat{\mathbf{x}},\theta).$ (18)
@@ 3.1 Bias identification with non-isotropic data densities
To illustrate the distinct methodologies, a modification of Plumlee’s “Pedagogical example"40 is implemented. In this case, the impact of non-isotropic data densities on the bias identification is evaluated. It presents a simple 1D case that is suitable for the different approaches. Taking $X=[0,1]$, the original measurements are generated from a "true" model governed by $y(x)=4x+x\sin 5x$ for $x\in X$, while the computer model is $f(x,\theta)=\theta x$. Measurements are collected for $\hat{X}=[0,0.25]\cup[0.4,0.5]\cup[0.8,1.0]$ at equidistant intervals of 0.05 and are perturbed with Gaussian noise of variance $\sigma^{2}=0.02$. The objective is to fit the variable $\theta$ such that the computer model reflects the measurements. There is a clear bias between $y$ and $f$, therefore it can potentially be quantified. Our modifications to the original example focus on highlighting the effect of an incomplete dataset and the comparison of the different approaches for the same system. To be used as a baseline model, the classical Bayesian approach without bias is implemented. Therefore, for a given $\theta$ and given evaluations $\hat{\bm{\epsilon}}$ of the Gaussian noise $\epsilon\sim\mathcal{N}(0.0,\sigma)$ for each observation point, the residuals of the predictions at the measurements can be expressed as
@@
The objective is to estimate a posterior distribution for $\theta$ using Bayesian inference. The prior for $\theta$ is modelled as a normal distribution $\theta\sim\mathcal{N}(2.0,0.5)$. A Gaussian model with prescribed additive error of $\sigma^{2}$ is defined as proposal distribution. The MCMC algorithm is run for 1000 steps with a burn-in of 100 steps, which is enough to ensure the chain convergence as shown in Table 1. The posterior predictive resulting from sampling the inferred $\theta$ is represented in Figure 1.
@@
It can be observed that the prediction has converged close to a Dirac’s delta distribution due to the measurement noise being prescribed with a small variance. The model cannot represent the measurements because of model discrepancy. Uncertainty is also underestimated due to such discrepancy. Additionally, the unbiased approach provides a value of $\theta$ that makes the predictions at the observation points converge to the minimum Mean Square Error (MSE) with respect to the measurements, with a value of 3.34 and a standard deviation of 0.01. However, that value does not coincide with the $L^{2}$ norm between the generating process and the computational model over the domain of interest $X=[0,1]$ due to undersampling of the regions $[0.25,0.4]$ and $[0.5,0.8]$. For reference, the optimal value of $\theta$ obtained using an $L^{2}$ loss is 3.6. This discrepancy further reduces the reliability of the inferred computational model far from the observation points, which is undesirable.
Next, the modularized version of KOH’s approach presented in Algorithm 1 is applied to the benchmark case. A GP with a Matérn kernel of $\nu=3/2$ is used to model the bias term $b$. The correlation length for the bias GP is fixed to $\ell_{b}=\frac{0.5}{\sqrt{3}}$ in the interest of a representative comparison between methods. In this case, the input is one-dimensional and requires only samples of $x$ for training, as the bias is defined for a given $\theta$ and will be newly trained at every MCMC evaluation. Consequently, the training set is generated at every sample of $\theta$ and is composed by the positions $\hat{\mathbf{x}}$ of the measurements as input, and their corresponding residuals $\hat{\mathbf{r}}$ for a given $\theta$ as output. Note that for this approach, the anchor points $\xi$ of the GP is always the input set $\hat{\mathbf{x}}$. The fitted parameters can be observed in Table 1, and its fitted and bias-corrected responses are represented in Figure 2. The bias is a GP regressor, therefore for the case of just one measurement per observation point and given a sufficient kernel, the bias-corrected predicted response interpolates the observations. However, in the regions without measurements, the response presents a significantly larger bias. The bias itself can be used as a quantification of the model uncertainty.
@@
Analogously to the other approaches, OGPs are implemented for the simple 1D case. The only difference compared to KOH’s is the imposition of the orthogonality condition between the bias and the latent parameters at every evaluation of the MCMC loop. The required derivatives are obtained by applying a simple finite differences method with a central scheme and an increment step $h=10^{-3}$. Every other parameter remains the same as for KOH, i.e. the prior distributions and correlation kernel. The inference results can be observed in Table 1, and Figure 3 present the predictive results. The chosen metric is the $L^{2}$ loss, therefore the expected optimal value for $\theta$ is 3.6.
@@
A comparison of the posterior predictive results is represented in Figure 4. It can be observed that none of the fitted models covers the data points without considering the bias. As expected, the no-bias and OGP approaches converge to a Dirac’s delta distribution, while KOH presents some variance. The analytical optimum value of $\theta$ using an $L^{2}$-norm without noise is 3.6, and for the Mean Square Error (MSE) on the predictions is 3.33. In the perfect case, the point estimate for the maximum likelihood of $\theta$ should converge to a chosen optimum with almost no variance in the parameter. Thanks to the definition of the anchor points $\hat{\mathbf{r}}$ over the whole domain, only OGP is able to achieve such precision and allows controlling the norm with respect to which the optimum is calculated.
@@ 3.2 Bias identification with heteroscedastic noise
This case aims to reflect the influence of model bias in the estimation of heteroscedastic noise. Two models are differentiated to this end: one simple cantilever beam under a deterministic load condition that is used as the prescribed computational model and another one biased and with a stochastic load used to generate the data. The objective is to infer the value of the Young’s modulus for the computational model based on observations obtained from the biased one. This simplification mimics the situation of a testing load on a concrete cantilever bridge. The computational model supposes that only the experimental load is present in the real bridge, while the observations may have been collected with additional loads on the bridge or even with traffic. Additionally, it is supposed that a deficient placement and setting of the sensors leads to an additional displacement of $\Delta u=-100$ mm for every measurement. These extreme additions are introduced for illustrative purposes, but comparable effects are observed regularly in real implementations of bridge monitoring systems, especially if the model updating is done with data from the bridge in service.
@@
Both models correspond to a FEM representation of such a bridge as a 2D cantilever beam of length $L=50$ m and height $H=3$ m, discretized with $30\times 10$ quadrilateral linear elements. This represents half of the span of a bridge with constant a cross-section. The measurements will predict the deformation of the beam under a load at the free end with a simple linear elastic material law. The load represents a typical static test truck of 100 kN. The assumed width of the bridge is $W=3$ and its Poisson’s ratio $\nu$ is 0.2. Young’s modulus $E$ is the subject of this study, but in general, it will have a base value of 30 GPa. A diagram for the model is represented in Figure 5.
@@
The complex generative model includes the additional prescribed displacement at the fixed boundary, simulating a displacement at the abutment or a deficient setting of the sensors. Additionally, in contrast with the deterministic load for the computational model of 100 kN, the generative model adds a stochastic component. The result is a constant force of 100 kN plus a random draw from a log-normal distribution with mean 20 kN and standard deviation 6 kN. This variability represents a simplification of the additional load from a vehicle over the bridge. Vertical displacement measurements are obtained at $n=5$ virtual sensor positions situated at $y=1.5$ m and $x=[10,20,30,40,50]$ m. The simulation is repeated 20 times, each with a different realization of the load, generating as many measurements per observation point. A prescribed Gaussian noise of zero mean and standard deviation of $2\times 10^{-8}$ m is added to the measurements to simulate typical sensor noise.
@@
In this case, the virtual sensors cover sufficiently well the observation domain, therefore the fitted latent parameters are similar with and without bias. However, two impactful sources of error have been introduced: the prescribed displacement at the boundary and the variation in the load. They create respectively a discrepancy in the model’s response compared to the measurements and additional noise with physical meaning that cannot be explained a priori. It is not possible to fully represent these effects without considering a bias term. First, the aforementioned discrepancy can be directly estimated by including the bias analogously as in the simple 1D case: adding it to the output as a GP with the same combination of constant and Matérn kernel. Second, to allow the same GP to consider multiple measurements per sensor coordinate, a noise-enabled kernel must be employed. Due to the noise coming from the variation in the load, its amplitude must reflect the physical behaviour of the beam, therefore the noise is expected to vary in scale at the different sensors. This requires employing a heteroscedastic noise component added to the kernel, whose parameters are trained during the phase where the bias is fitted. Such a component is using the heteroscedastic kernel implemented in gp_extras45. A fixed homoscedastic white noise component with a $\sigma$ value of 1e-12 m is added to avoid numerical inconsistencies due to the proximity to zero noise at the sensors near the cantilever boundary. The additional prescribed sensor noise is included in the regression kernel. Effectively, both the homoscedastic and heteroscedastic parts of the kernel will represent the same noise in the displacements produced by the load variability, which is expected to be larger than 1e-12 m at any sensor. Therefore, prescribing a constant homoscedastic part does not substantially impact the results, while increasing the numerical stability of the inference procedure. The additional derivatives required for OGP’s approach will be calculated using a central finite difference scheme with a step of $h=1$ GPa. An alternative popular approach for estimating this noise term in the absence of bias is to include it as a parameter to be inferred. This approach increments the number of latent parameters, does not address the model discrepancy and requires prior information on the noise that is not generally available. Nevertheless, it will be applied in the case with no bias in the interest of a more detailed comparison, providing further noise estimation capabilities.
@@
Convergence is achieved with every method in the chosen number of steps. For the approach without bias, the noise is inferred as an extra latent parameter. It must be noted $E$ and $\sigma$ do not allow for a unique identification of such parameters due to the effect of the model discrepancy. Nevertheless, the convergence is apparent. Independently of the scale of the Gaussian noise $\sigma$, the fitted model will not be able to replicate the generated dataset. In comparison with the real Young’s modulus with a mean of 30 GPa and a standard deviation of 10 GPa, the inference procedure converges to a mean of around 14.9 GPa and standard deviation 0.5 GPa, which means an error of more than 50%. This relatively large difference is expected, as the predictions will never be able to accommodate the observations due to the prescribed displacement in the generative process. To obtain predictions closer to the observations, the computational model tends to become less stiff to accommodate the offset in the observations. The magnitude of the inferred noise without bias remarkably ranges between 0.051 and 0.068 m. With these values, it would fail to represent the noise both near the fixed end, close to zero, and at the free end, of approximately 0.025 m for the sampled data. Therefore, these predictions from the fitted model do not guarantee that the computational model reflects the real system, which is a requirement in the implementation of digital twins.

### v2

@@ 2.3.1 Identification of noise in biased models
In general, defining the bias term as a GP process requires the selection of suitable kernels. Typical choices are Radial Basis Functions (RBF) or Matérn kernels49, aimed at regression problems. In the implementation of the GP, a prescribed white noise of variance $\sigma_{n}^{2}$ is usually added as in Equation 17. It provides numerical stability and, in some cases, represents the measurement error $\epsilon_{\text{meas}}$. Due to its flexible structure as a GP, the bias term will absorb $\epsilon_{\text{meas}}$, thus including it explicitly is not required. It is generally not possible to distinguish between $\epsilon_{\text{model}}$ and the non-prescribed $\epsilon_{\text{meas}}$. We will then consider any variation in the model discrepancy as part of $\epsilon_{\text{model}}$ and susceptible to be explained by the bias term. If more than one distinct output observation is available per input, the fitted GP tends to solve the regression problem and will not capture that variability. We consider these variations in the observations as noise to be identified, hence a noise-aware kernel with a non-vanishing variance at the observation points is required in those cases. This concept is analogous to the nugget effect considered in geostatistics51. In that context, we formalize a noise-aware kernel $k(x_{i},x_{j})=k(h)$ where $h$ is the lag or distance between $x_{i}$ and $x_{j}$, as such for which the semivariogram $\gamma(h)$ 52 of the resulting GP returns a non-zero value when $h\to 0$:
$\lim_{h\to 0}\gamma(h)\neq 0.$ (23)
Homoscedastic and heteroscedastic noise kernels are available to be added to the GP. Homoscedastic ones, such as white noise kernels, allow us to identify constant noise added to the observations. As the bias term aims to compensate for the discrepancy in the output of the model and $\epsilon_{\text{meas}}$ is assumed to be prescribed, homoscedastic kernels in the bias GP would capture additional observation noise of non-physical nature, such as deficient data collection. A typical homoscedastic kernel is a white noise kernel with noise variance $\sigma_{n}^{2}$ such that
@@
where $g(x_{i})$ represents the evaluation of the regressor built on the pairs $(\bm{\xi}_{i},\sigma_{i}^{2})$. Determining the set of anchor points $\hat{\bm{\xi}}$ is not straightforward in real applications. They control the behaviour of $g(x_{i})$: too many anchor points become expensive to compute and may lead to overfitting, while too few may not capture the variation in $\sigma_{i}^{2}$. It is general practice to take the training points of the GP as anchor points 53. If no additional information is available, this choice reflects the expectations that the training points will capture sufficiently well both bias and noise variance. Further considerations on how to model heteroscedasticity in GPs can be found in the work of Binois et al54.
In the implementation used in this work55, $g(x_{i})$ is obtained from applying kernel regression with the training points as anchor points $\hat{\bm{\xi}}$. This kernel can be added analogously to the homoscedastic one to the regressor kernel of the bias GP, where its contribution to the covariance matrix $C_{HN}$ is presented as a diagonal matrix with different entries obtained from the kernel evaluation:
@@
The introduction of noise-aware kernels in the bias allows quantifying the variability in the observations without modifying the deterministic computational model. These kernels can be implemented in both KOH and OGP approaches without needing to further modify the methodology presented in the previous sections. In the case of OGPs, the orthogonality condition is not affected by the noise introduction, as it is part of the kernel evaluation. Sung et al. 53 verified that the introduction of heteroscedastic noise with OGP bias improved the calibration of models in comparison with taking homoscedastic noise with no bias or KOH’s approach with no noise. In this paper, the comparison is extended to the framework of their applicability to digital twins of real structures.
@@ 2.3.2 Model enhancement through bias’ input domain extension
One of the most prominent advantages of implementing the bias term independently from the model is the possibility of including information in the bias that enriches and guides the model inference without needing to modify the computational model itself. Most approaches that extend a given computational model when bias is present use data-driven models that can be easily modified 6, 56, iteratively update the models 19, 34, 35, 36, or modify the model by defining functional outputs57. A review on such approaches can be found in Baker et al. 58 and Sung et al. 59 In contrast, the aspect presented here has the potential to enhance the predictions generated by simulation-based digital twins without modifying the model itself, only through a modification of the bias representation.
Being $f(x,\theta)$ the computational model, we can extend the model with a bias $b(x,\eta)$ such that $\eta$ are parameter coordinates not considered in $f$ but for which observations are available. Therefore, obviating the measurement noise $\epsilon_{\text{meas}}$, the modified Equation 2 for the final model would be
@@
which would be defined over the extended domain $(x,\eta)\in X\times H$. For example, a temperature-agnostic model defined over the spatial domain $X$ could be extended to consider the temperature by including it in the bias term. This approach requires the datasets used for calibration to be labelled for the extension variable as well as for the original ones. Such is the case for many digital twins, especially for large structures such as bridges, where abundant information from a variety of sources is gathered independently for monitoring purposes but not necessarily included in the model definition.
The introduction of an extension of the domain requires minor modifications to the inference workflow. In particular, the input training data is composed of pairs of $(x,\eta)$, while the computational model is only defined for $x$. Nevertheless, the GP for the bias takes as output training data the residuals, which are defined over $(x,\eta)$, which allows for keeping the associated pairs of $x$ and $\eta$. Therefore, for the training data of the GP, we define as input dataset the pairs $(\hat{\mathbf{x}},\hat{\bm{\eta}})$ and as output dataset the residuals for a given sampled $\bm{\theta}$
$\hat{\mathbf{r}}(\hat{\mathbf{x}},\hat{\bm{\eta}},\bm{\theta})=y(\hat{\mathbf{x}},\hat{\bm{\eta}})-f(\hat{\mathbf{x}},\bm{\theta}).$ (29)
@@ 3.1 Bias identification with non-isotropic data densities
To illustrate the distinct methodologies, a modification of Plumlee’s “Pedagogical example"42 is implemented. In this case, the impact of non-isotropic data densities, i.e. gaps in the measurements, on the bias identification is evaluated. Table 1 presents a simple 1D case that is suitable for the different approaches. Taking $X=[0,1]$ as the complete domain of interest, the original response is generated from a "true" model governed by $y(x)=4x+x\sin 5x$ for $x\in X$, while the computer model is $f(x,\theta)=\theta x$. Measurements are collected for $\hat{X}=[0,0.25]\cup[0.4,0.5]\cup[0.8,1.0]$ at equidistant intervals of 0.05 and are perturbed with Gaussian noise of variance $\sigma^{2}=0.02$. It is not uncommon that in practice measurements are not available for some parts of the domain of interest, represented in this case by the arbitrarily chosen intervals where no measurements are collected. This “gaps” in the measurement collection generates the aforementioned non-isotropic data densities, where some intervals of $X$ are represented with fewer measurements in the training set.
Element Formulation
Generator model $y(x)=4x+x\sin 5x$
Computational model without bias term $f(x,\theta)=\theta x+\epsilon$
Computational model with bias term $z(x,\theta)=f(x,\theta)+b(x)=\theta x+b(x)+\epsilon$
Discrepancy source $x\sin 5x$
Latent parameter and prior $\theta,\leavevmode\nobreak\ p(\theta)\sim\mathcal{N}(2.5,1.5)$
Domain of interest $x\in X=[0,1]$
Measurements’ domain $\hat{x}\in\hat{X}=[0,0.25]\cup[0.4,0.5]\cup[0.8,1.0]$
Noise $\epsilon\sim\mathcal{N}(0,\sqrt{0.02})$
Table 1: Case description for Section 3.1 (“Pedagogical example" model).
The objective is to fit the variable $\theta$ such that the computer model reflects the measurements. There is a clear discrepancy between $y$ and $f$, therefore it can potentially be quantified. Our modifications to the original example focus on highlighting the effect of an incomplete dataset and the comparison of the different approaches for the same system. To be used as a baseline model, the classical Bayesian approach without bias is implemented. Therefore, for a given $\theta$ and given evaluations $\hat{\bm{\epsilon}}$ of the Gaussian noise $\epsilon\sim\mathcal{N}(0.0,\sigma)$ for each observation point, the residuals of the predictions at the measurements can be expressed as
@@
The posterior distribution for $\theta$ is estimated using Bayesian inference. The prior for $\theta$ is modelled as a normal distribution $\theta\sim\mathcal{N}(2.5,1.5)$. This prior represents a poor initial belief with a high degree of uncertainty that still includes the observations in a reasonable range (see Figure 1). A Gaussian model with prescribed additive error of $\sigma^{2}$ is defined as proposal distribution. The MCMC algorithm is run for 1000 steps with a burn-in of 100 steps, which is enough to ensure the chain convergence as shown in Table 2. The posterior predictive resulting from sampling the inferred $\theta$ is represented in Figure 1.
@@
It can be observed that the prediction has converged close to a Dirac’s delta distribution due to the measurement noise being prescribed with a small variance. The model cannot represent the measurements because of model discrepancy. Uncertainty is also underestimated due to such discrepancy. Additionally, the unbiased approach provides a value of $\theta$ that makes the predictions at the observation points converge to the minimum Mean Square Error (MSE) with respect to the measurements, with a value of 3.34 and a standard deviation of 0.01. For reference, the optimal value of $\theta$ obtained using an $L^{2}$ loss (see Equation 5) is 3.6. The value obtained does not coincide with the $L^{2}$ norm between the generating process and the computational model over the domain of interest $X=[0,1]$ due to undersampling of the regions $[0.25,0.4]$ and $[0.5,0.8]$. This discrepancy further reduces the reliability of the inferred computational model far from the observation points, which is undesirable.
Next, the modularized version of KOH’s approach presented in Algorithm 1 is applied to the benchmark case. A GP with a Matérn kernel of $\nu=3/2$ is used to model the bias term $b$. The correlation length for the bias GP is fixed to $\ell_{b}=\frac{0.5}{\sqrt{3}}$ in the interest of a representative comparison between methods. In this case, the input is one-dimensional and requires only samples of $x$ for training, as the bias is defined for a given $\theta$ and will be newly trained at every MCMC evaluation. Consequently, the training set is generated at every sample of $\theta$ and is composed by the positions $\hat{\mathbf{x}}$ of the measurements as input, and their corresponding residuals $\hat{\mathbf{r}}$ for a given $\theta$ as output. Note that for this approach, the anchor points $\xi$ of the GP is always the input set $\hat{\mathbf{x}}$. The fitted parameters can be observed in Table 2, and its fitted and bias-corrected responses are represented in Figure 2. The bias is a GP regressor, therefore for the case of just one measurement per observation point and given a sufficient kernel, the bias-corrected predicted response interpolates the observations. However, in the regions without measurements, the response presents a significantly larger bias.
@@
Analogously to the other approaches, OGPs are implemented for the simple 1D case. The main difference compared to KOH’s is the imposition of the orthogonality condition between the bias and the latent parameters at every evaluation of the MCMC loop. The required derivatives are obtained by applying a simple finite differences method with a central scheme and an increment step $h=10^{-3}$. Additionally, the anchor points $\bm{\xi}$ are defined over the whole input domain $X$ discretized by the equidistant points $\mathbf{x}$ separated by intervals of 0.05. Consequently, the anchor points prescribe the orthogonality condition as well on the regions of the domain of interest without measurements. Every other parameter remains the same as for KOH, i.e. the prior distributions and correlation kernel. The inference results can be observed in Table 2, and Figure 3 present the predictive results. The chosen metric is the $L^{2}$ loss, therefore the expected optimal value for $\theta$ is 3.6.
@@
A comparison of the posterior predictive results is represented in Figure 4. It can be observed that none of the fitted models covers the data points without considering the bias. As expected, the no-bias and OGP approaches converge to a Dirac’s delta distribution, while KOH presents some variance. The analytical optimum value of $\theta$ using an $L^{2}$-norm without noise is 3.6, and for the Mean Square Error (MSE) on the predictions is 3.33. In the perfect case, the point estimate for the maximum likelihood of $\theta$ should converge to a chosen optimum with almost no variance in the parameter. Thanks to the definition of the anchor points $\hat{\bm{\xi}}$ over the whole domain, only OGP is able to achieve such precision and allows controlling the norm with respect to which the optimum is calculated.
@@ 3.2 Bias identification with heteroscedastic noise
This case aims to reflect the influence of model bias in the estimation of heteroscedastic noise. Two models are differentiated to this end: one simple cantilever beam under a deterministic load condition that is used as the prescribed computational model and another one biased and with a stochastic load used to generate the data. The objective is to infer the value of the Young’s modulus for the computational model based on observations obtained from the biased one. This simplification mimics the situation of a testing load on a concrete cantilever bridge. The computational model supposes that only the experimental load is present in the real bridge, while the observations may have been collected with additional loads on the bridge or even with traffic. Additionally, it is supposed that a deficient placement and setting of the sensors leads to an additional displacement of $\Delta u=-100$ mm for every measurement. These extreme additions are introduced for illustrative purposes, but comparable effects are observed regularly in real implementations of bridge monitoring systems, especially if the model updating is done with data from the bridge in service. Table 4 summarizes the case’s formulation.
Element Formulation
Generator model assumptions Linear elastic cantilever beam that depends on the Young’s modulus $E$ and the loads vector $\bm{\phi}$, and outputs the vertical displacements vector $\bm{u}$ at $\bm{x}$. $\Delta u(x=0\text{ m})=-0.1\text{ m}$, $\phi(x=50\text{ m})=100\text{ kN}+\delta_{\phi}$ with $\delta_{\phi}\sim\mathcal{LN}(20\text{ kN},6\text{ kN})$.
Generator model $u^{g}(x,E)=g(x,E)+\epsilon$ for a given realization of $\delta_{\phi}$
Computational model assumptions Linear elastic cantilever beam that depends on the Young’s modulus $E$ and the loads vector $\bm{\phi}$, and outputs the vertical displacements vector $\bm{u}$ at $\bm{x}$. $\Delta u(x=0\text{ m})=0.0\text{ m}$, $\phi(x=50\text{ m})=100\text{ kN}$.
Computational model without bias term $u^{f}(x,E)=f(x,E)+\epsilon$
Computational model with bias term $u^{h}(x,E)=u^{f}(x,\theta)+b(x)=f(x,E)+b(x)+\epsilon$
Discrepancy source $\Delta u,\leavevmode\nobreak\ \delta_{\phi}$
Latent parameter and prior $E,\leavevmode\nobreak\ p(E)\sim\mathcal{N}(35\text{ GPa},5\text{ GPa})$, (with no bias, additionally $\sigma_{\epsilon},\leavevmode\nobreak\ p(\sigma_{\epsilon})\sim\mathcal{U}(0,0.8)$)
Domain of interest $x\in X=[10\text{ m},50\text{ m}]$
Measurements’ domain $\hat{x}\in\hat{X}=[10\text{ m},50\text{ m}]$, observations at $\hat{x}=[10\text{ m},20\text{ m},30\text{ m},40\text{ m},50\text{ m}]$
Noise $\epsilon\sim\mathcal{N}(0,\sqrt{2\times 10^{-8}})$
Table 4: Case description for Section 3.2 (beam model). The terms $f(x,E)$ and $g(x,E)$ represent the output of the FE model of the system under the assumptions of the computational model and the generator, respectively.
@@
Both models correspond to a Finite Elements (FE) 61 representation of such a bridge as a 2D cantilever beam of length $L=50$ m and height $H=3$ m, discretized with $30\times 10$ quadrilateral linear elements. This represents half of the span of a bridge with constant a cross-section. The measurements will predict the deformation of the beam under a load at the free end with a simple linear elastic material law. The load represents a typical static test truck of 100 kN. The assumed width of the bridge is $W=3$ and its Poisson’s ratio $\nu$ is 0.2. Young’s modulus $E$ is the subject of this study, but in general, it will have a base value of 30 GPa. A diagram for the model is represented in Figure 5.
@@
The complex generative model includes the additional prescribed displacement at the fixed boundary, simulating a displacement at the abutment or a deficient setting of the sensors. Additionally, in contrast with the deterministic load for the computational model of 100 kN, the generative model adds a stochastic component. The result is a constant force of 100 kN plus a random draw from a log-normal distribution with mean 20 kN and standard deviation 6 kN. This variability represents a simplification of the additional load from a vehicle over the bridge. The origin of coordinates is located at the bottom part of the fix end of the beam. Vertical displacement measurements are obtained at $n=5$ virtual sensor positions situated at $\hat{y}=1.5$ m and $\hat{x}=[10\text{ m},20\text{ m},30\text{ m},40\text{ m},50\text{ m}]$, which correspond to points in the middle line of the beam. The simulation is repeated 20 times, each with a different realization of the load, generating as many measurements per observation point. A prescribed Gaussian noise of zero mean and standard deviation of $2\times 10^{-8}$ m is added to the measurements to simulate typical sensor noise.
@@
In this case, the virtual sensors cover sufficiently well the observation domain, therefore the fitted latent parameters are similar with and without bias. However, two impactful sources of error have been introduced: the prescribed displacement at the boundary and the variation in the load. They create respectively a discrepancy in the model’s response compared to the measurements and additional noise with physical meaning that cannot be explained a priori. It is not possible to fully represent these effects without considering a bias term. First, the aforementioned discrepancy can be directly estimated by including the bias analogously as in the simple 1D case: adding it to the output as a GP with the same combination of constant and Matérn kernel $k_{GP}$. Second, to allow the same GP to consider multiple measurements per sensor coordinate, a noise-enabled kernel must be employed. Due to the noise coming from the variation in the load, its amplitude must reflect the physical behaviour of the beam, therefore the noise is expected to vary in scale at the different sensors. This requires employing a heteroscedastic noise component added to the kernel, whose parameters are trained during the phase where the bias is fitted. Such a component is using the heteroscedastic kernel implemented in gp_extras55. A fixed homoscedastic white noise component with a $\sigma_{N}$ value of 1e-12 m is added to avoid numerical inconsistencies due to the proximity to zero noise at the sensors near the cantilever boundary. The resulting kernel follows
$k(x_{i},x_{j})=k_{GP}(x_{i},x_{j})k_{WN}(x_{i},x_{j})k_{HN}(x_{i},x_{j}),$ (31)
where $k_{WN}(x_{i},x_{j})$ is defined as in Equation 24 and $k_{HN}(x_{i},x_{j})$ as in Equation 26 where $\bm{\xi}=\hat{\mathbf{x}}$ and the $\sigma_{i}$ are fitted during the model evaluation.
Effectively, both the homoscedastic and heteroscedastic parts of the kernel will represent the same noise in the displacements produced by the load variability, which is expected to be larger than 1e-12 m at any sensor. Therefore, prescribing a constant homoscedastic part does not substantially impact the results, while increasing the numerical stability of the inference procedure. The additional derivatives required for OGP’s approach will be calculated using a central finite difference scheme with a step of $h=1$ GPa. An alternative popular approach for estimating this noise term in the absence of bias is to include it as a parameter to be inferred. This approach increments the number of latent parameters, does not address the model discrepancy and requires prior information on the noise that is not generally available. Nevertheless, it will be applied in the case with no bias in the interest of a more detailed comparison, providing further noise estimation capabilities.
@@
Convergence is achieved with every method in the chosen number of steps. For the approach without bias, the noise is inferred as an extra latent parameter. It must be noted $E$ and $\sigma_{\epsilon}$ do not allow for a unique identification of such parameters due to the effect of the model discrepancy. Nevertheless, the convergence is apparent. Independently of the scale of the Gaussian noise $\sigma$, the fitted model will not be able to replicate the generated dataset. In comparison with the real Young’s modulus with a mean of 30 GPa and a standard deviation of 10 GPa, the inference procedure converges to a mean of around 14.9 GPa and standard deviation 0.5 GPa, which means an error of more than 50%. This relatively large difference is expected, as the predictions will never be able to accommodate the observations due to the prescribed displacement in the generative process. To obtain predictions closer to the observations, the computational model tends to become less stiff to accommodate the offset in the observations. The magnitude of the inferred noise without bias remarkably ranges between 0.051 and 0.068 m. With these values, it would fail to represent the noise both near the fixed end, close to zero, and at the free end, of approximately 0.025 m for the sampled data. Therefore, these predictions from the fitted model do not guarantee that the computational model reflects the real system, which is a requirement in the implementation of digital twins.